OpenStudy (anonymous):

You are dealt one card from a 52-card deck. Find the probability that you are not dealt a spade. A 4/13 B 1/4 C 2/5 D 3/4

OpenStudy (anonymous):

numbered card, that is 2 through 10 or 9 per suite or 36 total, so the probability is 36/52 = 9/13 (if you count the ace as a numbered card, that changes it) club, 1/4 If that means OR, or the sum of the two, that is more complicated because of overlap. clubs are still 1/4, and number card for the other 3 suites is 9/39 = 3/13 so overall it is 1/4+3/13 or (13+12)/52 = 25/52

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

Thanks !